OpenStudy (anonymous):

Rewrite this statement using brackets to make it true. 18+6/2x3

OpenStudy (anonymous):

18+6/2x(3)

OpenStudy (igreen):

Is this it? \(\dfrac{18+6}{2 \times 3}\)

OpenStudy (igreen):

@XavierEripmav

OpenStudy (anonymous):

Using brackets :/

OpenStudy (igreen):

Well, if that is it, we can put brackets around the multiplication..because in PEMDAS, multiplication comes BEFORE addition.

OpenStudy (igreen):

So just put a bracket around 2 x 3.

OpenStudy (igreen):

That gives you: \(\dfrac{18+6}{[2 \times 3]}\)

OpenStudy (igreen):

Remember, Brackets are the FIRST step in PEMDAS. So to keep this the same, we have to put the brackets around the multiplication, because multiplication comes before addition.

OpenStudy (anonymous):

Your right, thanks a lot.

OpenStudy (igreen):

No problem.

OpenStudy (igreen):

Actually, it really wouldn't matter if you added first in this situation, but I'm pretty sure that's what they want you to do.